OpenStudy (anonymous):

Can anyone explain to me how to simplify an expression with negative exponents using the rules for multiplying monomials? Have to know it and im not sure!!

OpenStudy (ranga):

Not sure if this is what you are looking for: \[\Large a^{-n} = \frac{ 1 }{ a^n }\]

OpenStudy (ranga):

You can give an example from the book and we can try simplifying it.

OpenStudy (anonymous):

I am online school, and super confused!!! It doesnt give an example, and honestly, I'm not sure if that is what I'm lookin for or not.

OpenStudy (ranga):

There are a few rules involving exponents. Since you were specifically asking about negative exponents I thought you can use the above identity and convert a negative exponent into a positive exponent and use the rules of the exponents just the way you would with a positive exponent.

OpenStudy (anonymous):

Ok? I am so sorry, I super confused. :/

OpenStudy (ranga):

Example: \[\Large y^5 * y^{-8} = y^{5 + (-8))} = y^{-3} = \frac{ 1 }{ y^3 }\]

OpenStudy (ranga):

The expression on the left has positive and negative exponents. To simplify it I used a couple of rules of exponents and simplified it.

OpenStudy (ranga):

Here are the exponent rules: http://www.rapidtables.com/math/number/exponent.htm

OpenStudy (ranga):

Okay, I understand the question now. Multiplying two monomials and simplifying them is what you are looking for except there are negative exponents involved. The example I gave before is one. Here is another example: Multiply and simplify:\[\Large x^{-8} * x^{-5} = x^{-8 + (-5))} = x^{-8-5} = x^{-13} = \frac{ 1 }{ x^{13} }\]

OpenStudy (anonymous):

Ok thanks!!

OpenStudy (ranga):

you are welcome.